# Aspects of the geometry and topology of expanding horizons

### v1

@@ 2. Preliminaries
Then the null expansion $\theta^{+}$ of $\Sigma^{\prime}$ cannot satisfy $\theta^{+}<\theta_{0}$. This follows from e.g. [15, Prop. 2], which otherwise implies, under the condition (2.6), that there is a null focal point along any such $\eta$. In the case when matter (e.g. an accretion disk) is present in the vicinity of $\Sigma$ (interpreted as a cross-section of the event horizon with small but positive outward null expansion), this provides some justification for the assumption that $\Sigma$ is $\mathfrak{g}$-locally weakly outermost, and hence $\mathfrak{g}$-stable.
3. Proofs
3.1. Proof of Theorem 1.1
@@ 3. Proofs
$\mathcal{P}=\mu+J(\nu)+\frac{1}{2}|\chi^{+}|^{2}+\frac{1}{2}h^{2}-h\tau\geq\mu+J(\nu)-h\tau\geq 0,$
where the last inequality follows from either of the assumptions (i) and (ii), since $\mu+J(\nu)\geq\mu-|J|$. If $P\equiv 0$, then the above inequalities imply that $h\tau=\mu+J(\nu)$, which in turn implies $h\equiv 0$. This contradicts both (i) and (ii), since in the latter case we have that $h\tau=c_{0}>0$. The result then follows from Lemma 2.3
3.2. Proof of Theorem 1.2
First, observe that it is not difficult to construct a spacelike hypersurface $\hat{M}$ in $(\bar{M},\bar{g})$, with $\Sigma\subset\hat{M}$, satisfying the following conditions (see Figure 1):
  • $\hat{M}$ meets $M$ tangentially along $\Sigma$;
  • $\hat{M}$ is in the causal past of $M$;
  • $\hat{M}$ has mean curvature $\hat{\tau}\leq 0$.
If $\Sigma$ is $\mathfrak{g}$-stable in $\hat{M}$, then we can apply Theorem 1.1 to conclude that $\Sigma$ admits a metric of positive scalar curvature.
Suppose, by contradiction, that $\Sigma$ is not $\mathfrak{g}$-stable in $\hat{M}$. In particular, for each outward neighborhood $\hat{U}_{0}$ of $\Sigma$ in $\hat{M}$, there exists a hypersurface $\hat{\Sigma}$ whose null expansion $\hat{\theta}^{+}$ satisfies $\hat{\theta}^{+}<\theta_{0}$ everywhere. Now, let $\mathcal{H}$ be the null hypersurface generated by the future directed outward null geodesics orthogonal to $\hat{\Sigma}$. Taking a smaller outward neighborhood $\hat{U}_{0}$ if necessary, $\mathcal{H}$ is a smooth null hypersurface that meets $M$ in a closed hypersurface $\tilde{\Sigma}$ contained in $U_{0}$. Then, the Raychaudhuri equation for null geodesic congruences together with the null energy condition (which is a consequence of the dominant energy condition) give that the null expansion $\tilde{\theta}^{+}$ of $\tilde{\Sigma}$ is (pointwise) less than or equal to $\hat{\theta}^{+}$; hence, $\tilde{\theta}^{+}<\theta_{0}$ everywhere, which contradicts the assumption that $\Sigma$ is $\mathfrak{g}$-locally weakly outermost in $M$.
Figure 1.
3.3. Proof of Theorem 1.3
@@
Above, we have used the dominant energy condition. Since $h\theta^{-}\not\equiv 0$, the result follows from Lemma 2.3.
3.4. Proof of Theorem 1.4
@@
Consider the modified initial data set $(M,g,\hat{K})$, where $\hat{K}=K-\frac{h}{n-1}g$. On $U$, the function $h$ is defined by
$h(t)=\begin{cases}\theta_{0}\exp\left(1-\frac{1}{1-(\frac{t}{\epsilon})^{2}}\right),&0\leq t<\epsilon,\\ 0,&\epsilon\leq t\leq 2\epsilon,\end{cases}$
@@
Note that $\hat{\theta}^{+}\coloneqq\operatorname{tr}_{\Sigma}(\hat{K}+A)=\theta^{+}-\theta_{0}\leq 0$, where $A$ is the second fundamental form of $\Sigma$ in $(M,g)$ with respect to the normal pointing into $M$; that is, $\Sigma$ is weakly outer trapped in $(M,g,\hat{K})$. Further, the DEC clearly holds on $M\setminus U$. On $U$, a computation shows that (cf. [19, Section 6])
@@
The theorem then follows from [18, Theorem 1.3].

### v2

@@ 2.2. Stability with respect to null hypersurfaces
    for all $\psi\in C^{\infty}(\Sigma)$. Hence, the first eigenvalue $\lambda_{1}(L_{0})$ of $L_{0}=-\Delta+\mathcal{Q}$ is nonnegative. Furthermore, if equality in (2.4) holds, and $\psi\not\equiv 0$, then $\psi$ is an eigenfunction of $L_{0}$ associated with $\lambda_{1}(L_{0})=0$;
  2. (ii)
    If $\mathcal{Q}=\frac{1}{2}R_{\gamma}-\mathcal{P}$ for some function $\mathcal{P}\geq 0$, then $\Sigma$ admits a metric of positive scalar curvature, unless $\lambda_{1}(L)=0$, $\mathcal{P}\equiv 0$, and $(\Sigma,\gamma)$ is Ricci flat.
2.3. Further comments on the $\mathfrak{g}$-stability condition
As briefly noted in the introduction, under suitable conditions, cross sections $\Sigma$ of the event horizon in stationary black hole spacetimes are MOTS (this follows from, for example, results in [9]). Moreover, under general circumstances, such a MOTS $\Sigma$ will be stable. For suppose not. Then $\Sigma$ could be pushed outside the event horizon along a spacelike (or null) hypersurface to an outer trapped surface $\hat{\Sigma}$. Now, it is a basic result in the theory of black holes that, for suitably defined black hole spacetimes satisfying the null energy condition, ${\rm Ric}(X,X)\geq 0$ for all null vectors $X$ (in particular, satisfying the vacuum Einstein equations, ${\rm Ric}\equiv 0$), trapped surfaces cannot exist outside the event horizon; see e.g. [26, Prop. 12.2.2]. The proof is easily modified to also rule out outer trapped surfaces that are homologous to the cross section $\Sigma$, such as $\hat{\Sigma}$.
The point we wish to make here is that essentially the same conclusion of stability can be reached for cross sections $\Sigma$ of expanding horizons, provided there is sufficient matter in the vicinity of the black hole.
Let $\Sigma$ be a cross section with null expansion $h$ bounded above by a constant $\theta_{0}>0$. If $\Sigma$ is not stable, then it can be pushed outside the event horizon along a spacelike (or null) hypersurface to a surface $\hat{\Sigma}$ having null expansion $\hat{h}<\theta_{0}$. Suppose however that, due to the presence of matter, along each future complete outward null normal geodesic $s\longmapsto\eta(s)$ to $\hat{\Sigma}$, with $\eta^{\prime}(0)=\ell^{+}(p)$, $p=\eta(0)\in\hat{\Sigma}$,
[…]
Then the null expansion $\hat{h}$ of $\hat{\Sigma}$ cannot satisfy $\hat{h}<\theta_{0}$, for if it did then e.g. [17, Prop. 2] would imply that there is a null focal point along every such $\eta$. However, arguing as in the proof of [26, Prop. 12.2.2], there must exist at least one $\eta$ without a focal point.
3. Properties expanding horizons
Here we present the results briefly mentioned in the introduction.
3.1. The topology of expanding horizons
Theorem 3.1.
Let $(M,g,K)$ be an initial data set and $\Sigma$ be a closed hypersurface in $M$. If $\Sigma$ is a $\mathfrak{g}$-stable surface in $(M,g,K)$ with null expansion $\theta^{+}=h\in C^{\infty}(\Sigma)$, $h\geq 0$, and either
  1. (i)
    $\mu-|J|\geq 0$ and $\tau\leq\frac{1}{2}h$ along $\Sigma$, $h\not\equiv 0$, or
  2. (ii)
then $\Sigma$ admits a metric of positive scalar curvature.
As is well known, there are many restrictions on the topology of manifolds that admit a metric of positive scalar curvature (see e.g. [8] for a recent reference). We note that the condition on $\tau$ (which, recall, is the trace of $K$) in case (i) is satisfied if $\tau\leq 0$, and hence, in particular, if $M$ is maximal, i.e. $\tau=0$. It is also satisfied, of course, if $h$ has a positive lower bound $h_{0}$ and $\tau\leq\frac{1}{2}h_{0}$. A result related to Theorem 3.1 was obtained in [7] using a slightly different version of stability, one which is well-adapted to constructing surfaces of prescribed null expansion.
Proof.
@@ 3.1. The topology of expanding horizons
$\mathcal{P}=\mu+J(\nu)+\frac{1}{2}|\chi^{+}|^{2}+\frac{1}{2}h^{2}-h\tau.$ (3.1)
Case (i): In this case, since $\mu+J(\nu)\geq\mu-|J|\geq 0$, we have
$\mathcal{P}\geq\mu+J(\nu)+h\left(\frac{1}{2}h-\tau\right)\geq 0$
by the assumptions on $h$ and $\tau$. If $P\equiv 0$, then the above inequalities give that
$h\left(\frac{1}{2}h-\tau\right)=-(\mu+J(\nu)).$
Substitution of this into (3.1) then implies $\chi^{+}\equiv 0$, and hence $h=\operatorname{tr}\chi^{+}\equiv 0$, which is a contradiction. Thus, $P\not\equiv 0$.
Case (ii): In this case,
$\mathcal{P}\geq\mu+J(\nu)-h\tau\geq 0,$
since $\mu+J(\nu)\geq c_{0}$ and $h\tau\leq c_{0}$. If $P\equiv 0$, then the above inequalities imply $h\tau=\mu+J(\nu)$. Substitution of this into (3.1) gives that $h\equiv 0$. On the other hand,
$c_{0}\geq h\tau=\mu+J(\nu)\geq c_{0},$
and hence $h\tau=c_{0}>0$, which is a contradiction. Thus, $P\not\equiv 0$.
In either case, the result now follows from Theorem 2.5.∎
The next result is a spacetime result, in which $\Sigma$ is embedded in a null hypersurface. Physically, we may think of $\Sigma$ as arising from the transverse intersection of the event horizon (assumed to be sufficiently smooth) with this hypersurface. The theorem assumes the (spacetime) dominant energy condition, and a natural condition on the inner null expansion of $\Sigma$.
Theorem 3.2.
Proof.
@@
Above, we have used the dominant energy condition. Since $h\theta^{-}\not\equiv 0$, the result follows from Theorem 2.5.∎
3.2. Area bounds for expanding horizons
Theorem 3.3.
Let $(M,g,K)$ be a $3$-dimensional initial data set and $\Sigma$ be a closed surface in $M$. Suppose $\Sigma$ is a $\mathfrak{g}$-stable surface in $(M,g,K)$ with null expansion $\theta^{+}=h\in C^{\infty}(\Sigma)$, $h\geq 0$, such that $\tau\leq\frac{1}{2}h$. If $\mu-|J|\geq c$ for some constant $c>0$, then $\Sigma$ is topologically a $2$-sphere and the area of $\Sigma$ satisfies,
$A(\Sigma)\leq\frac{4\pi}{c}.$ (3.2)
Furthermore, if $A(\Sigma)=4\pi/c$, then the following hold along $\Sigma$:
  1. (1)
    $\chi^{+}=0$ and, in particular, $\Sigma$ is a MOTS;
  2. (2)
    $\Sigma$ is a round $2$-sphere of constant Gaussian curvature $\kappa_{\Sigma}=c$;
  3. (3)
    $\mu+J(\nu)=\mu-|J|=c$ on $\Sigma$.
Remark 3.4.
Rigidity results deriving from equality in (3.2) have been examined in detail by the authors in [14] and [15]. We note, however, that equality is not realized for expanding horizons, $\theta^{+}=h>0$.
Proof.
It follows from (2.2) that
where
$\mathcal{Q}=\kappa_{\Sigma}-\mathcal{P}$
and
$\mathcal{P}$ $=\mu+J(\nu)+\frac{1}{2}|\chi^{+}|^{2}+\frac{1}{2}h^{2}-h\tau$
$\geq\mu-|J|+h\left(\frac{1}{2}h-\tau\right)\geq c.$
Hence, setting $\psi=1$ in Theorem 2.5 (i), we obtain,
[…]
where $g(\Sigma)$ is the genus of $\Sigma$. These inequalities force $g(\Sigma)=0$, and the area bound follows.
If $A(\Sigma)=4\pi/c$, then all above inequalities must be equalities. In particular,
  • $\chi^{+}=0$;
  • $\mu+J(\nu)=\mu-|J|=c$ on $\Sigma$;
  • $\psi=1$ is an eigenfunction of $L_{0}=-\Delta+\mathcal{Q}$ associated with the first eigenvalue $\lambda_{1}(L_{0})=0$, that is,
    $\mathcal{Q}=\kappa_{\Sigma}-\mathcal{P}=0\quad\therefore\quad\kappa_{\Sigma}=\mathcal{P}=c.$
Similarly, we have:
Theorem 3.5.
Let $(\bar{M},\bar{g})$ be a $(3+1)$-dimensional spacetime and $\mathcal{N}$ be a null hypersurface in $(\bar{M},\bar{g})$. Suppose $\Sigma$ is a $\mathfrak{g}$-stable surface in $\mathcal{N}$ with null expansion scalars $\theta^{+}=h\geq 0$ and $\theta^{-}<0$. If $G(\ell^{+},\ell^{-})\geq c$ for some constant $c>0$, then $\Sigma$ is topologically a $2$-sphere and the area of $\Sigma$ satisfies,
$A(\Sigma)\leq\frac{4\pi}{c}.$
Furthermore, if $A(\Sigma)=4\pi/c$, then the following hold along $\Sigma$:
  1. (1)
    $\Sigma$ is a MOTS;
  2. (2)
    $\Sigma$ is a round $2$-sphere of constant Gaussian curvature $\kappa_{\Sigma}=c$;
  3. (3)
    $G(\ell^{+},\ell^{-})=c$ on $\Sigma$.
Proof.
It follows from (2.3) that
where
$\mathcal{Q}=\kappa_{\Sigma}-\mathcal{P}$
and
$\mathcal{P}=G(\ell^{+},\ell^{-})-h\theta^{-}\geq c.$
Inequalities (3.3) again apply.
If $A(\Sigma)=4\pi/c$, then $\mathcal{P}=G(\ell^{+},\ell^{-})=c$ and $-h\theta^{-}=0$ on $\Sigma$; in particular, $\Sigma$ is a MOTS, since $\theta^{-}<0$. Also, as before, $\mathcal{Q}=0$, that is, $\kappa_{\Sigma}=\mathcal{P}=c$. ∎
Theorem 3.5 improves certain aspects of results in [19].
3.3. A positive mass theorem for expanding horizons
The next theorem pertains to the spacetime positive mass theorem [10]. Recent results have extended this theorem to initial data sets with boundary [13, 21], where the components are assumed to be weakly outer trapped ($\theta^{+}\leq 0$), or weakly inner untrapped ($\theta^{-}\geq 0$). Here we show that the weakly outer trapped condition can be relaxed a bit, by requiring the dominant energy condition (for initial data sets) to be sufficiently strict near $\Sigma$.
Theorem 3.6.
Proof.
@@ 3.3. A positive mass theorem for expanding horizons
Consider the modified initial data set $(M,g,\hat{K})$, where $\hat{K}=K-\frac{h}{n-1}g$. On $U$, the function $h$ is defined by,
$h(t)=\begin{cases}\theta_{0}\exp\left(1-\dfrac{1}{1-(\frac{t}{\epsilon})^{2}}\right),&0\leq t<\epsilon,\\ 0,&\epsilon\leq t\leq 2\epsilon,\end{cases}$
@@
Note that $\hat{\theta}^{+}\coloneqq\operatorname{tr}_{\Sigma}(\hat{K}+A)=\theta^{+}-\theta_{0}\leq 0$, where $A$ is the second fundamental form of $\Sigma$ in $(M,g)$ with respect to the normal pointing into $M$; that is, $\Sigma$ is weakly outer trapped in $(M,g,\hat{K})$. Further, the DEC clearly holds on $M\setminus U$. On $U$, a computation shows that (cf. [22, Section 6])
@@
The theorem then follows from [21, Theorem 1.3].